the ratio of the measures of two supplementary angles is 5:3. Find the measures of the angles.
Answer by checkley77(12569)

5:3=X:(180-X)
3X=5(180-X)
3X=900-5X
3X+5X=900
8X=900
X=900/8
X-112.5 ANS.
180-112.5=67.5 ANS.
-------------------------------------
OR:
THERE ARE 8 PARTS (5+3=8) WITHIN THE 180 DEGREES.
180/8=22.5 IS THE VALUE OF EACH PART.
3*22.5=67.5 DEGREES FOR THE SMALLER ANGLE.
5*22.5=112.5 DEGREES FOR THE LARGER ANGLE.
-------------------------------------------
PROOF:
67.5+112.5=180
180=180